OpenStudy (anonymous):

A boat has a speed of 15 mph in calm water. It takes the boat 3 hours to travel upstream but only 2 hours to travel the same distance downstream. Which equation can be used to find c, the speed of the current

OpenStudy (anonymous):

_____________ speed__________time__________distance Upstream_______ 15-c___________3_____________d=(15-c)*3{rate*time} Down stream_____15+c__________2_____________d=(15+c)*2 d is equal for both directions, so equate the expressions: (15-c)*3=(15+c)*2

OpenStudy (anonymous):

simplify it to get c..speed of current

OpenStudy (anonymous):

and hope you know why upstream speed is 15-c and down stream speed is 15+c

OpenStudy (anonymous):

simplification... 45-3c=30+2c 15=5c c=3mph